sqrt y-5 + sqrt y = 5
Answer by edjones(8007)

sqrt(y-5)+sqrt(y)=5
2sqrt(y^2-5y)+2y-5=25 square each side
2sqrt(y^2-5y)=30-2y
4(y^2-5y)=4y^2-120y+900 square each side
4y^2-20y=4y^2-120y+900
100y=900
y=9
.
Ed